Ask your own question, for FREE!
Mathematics 5 Online
OpenStudy (anonymous):

A package contains 18 candy canes, 13 of which are cracked. If 4 are selected, find the probability of getting exactly four cracked candy canes.

OpenStudy (anonymous):

Ok so the probability that the candy is cracked is \[P(C)= \frac{13}{18}=.722\] The Probability that a candy isnt cracked is P(W)=1-.722=.278 Are you familiar with the binomial distribution?

OpenStudy (anonymous):

Ok so basically this is the formula for the binomial distribution|dw:1384716610738:dw|

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
17 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
20 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!